OpenStudy (anonymous):

When a person is walking, each hip now supports up to 2.5 times his upper weight, depending on how fast he is walking. (Recall that when you walk, your weight shifts from one leg to the other and your body comes down fairly hard on each leg.) For a 65 kg person, what is the maximum kinetic friction force at the hip joint if Fk has its minimum value of 0.0050?

OpenStudy (ganpat):

Frictional force = Fk N = 0.0050 * mg = 0.0050 * 65 * 9.81= 3.18 N !! (Not sure ).. @Vaidehi09 ??

OpenStudy (anonymous):

I don't think Fk here stands for coefficient of kinetic friction, does it? since coeff of kinetic friction is a constant. so it cant have max/min values. I'd rather say Fk stands for kinetic friction force. do correct me if i'm wrong. @Ganpat sry, i'm also not really sure about this one.

OpenStudy (anonymous):

i'm not sure, but i got the answer as 0.0125 N.

OpenStudy (ganpat):

I am not very good at this, but when i thought, i got following reason.. The Fk has minimum value may be because of varying fluid inside the body.. The more fluid will reduce friction, while less body fluid will increase the friction co-efficient.. So the lubricating effect can be taken care of .. what say ??

OpenStudy (anonymous):

from what i gather, the varying value of Fk is because of this statement in the ques: "each hip now supports up to 2.5 times his upper weight, depending on how fast he is walking." notice the "upto".....so that means the variable here is x. whose range is from an initial value to 2.5. so Fk is max when x = 2.5. I'm not sure abt the initial value....i took it as 1.

OpenStudy (ganpat):

No Idea.. need to brush up my knowledge again.. its been long.. Sorry dude !!